OpenStudy (ledah):

The area of a triangle is 25 in² and the height of the triangle is 10 in. How long is the base of the triangle?

OpenStudy (anonymous):

\[\large Area(Triangle) = \frac{1}{2} \times Base \times Height\] Use this..

OpenStudy (ledah):

Thank you